# DualXDA: Towards Sparse, Efficient and Explainable Data Attribution in Large AI Models

### v2

@@ Appendix H Unifying Concept-Level Explanations, Feature Attribution and Data Attribution
The formulation of DualDA can be combined with concept-level explanations to understand the impact that individual concepts – which may be present or absent in the training and test data – have on their combined DualDA attribution value. For a given concept , assume through a concept-discovery method such as linear probes we have obtained a Concept Activation Vector (CAV) (Kim et al., 2018). The larger the dot product between the hidden activations and the vector , the more likely it is that the concept is present in the test sample. We further normalize the CAV . Then, the amount of the DualDA attribution that is explained by concept can be quantified as
@@
An illustrative application of this methodology to an ImageNet sample is presented in Figure H.1. The attribution decomposition is performed with respect to concepts encoded by neurons in the penultimate network layer. Heatmaps are visualized for the three neurons contributing the highest relevance scores. To determine their corresponding conceptual representations, we present the nine training samples that maximize the activation of each neuron: the first neuron responds predominantly to turtle imagery, the second exhibits selectivity for hemispherical structures oriented upward such as shells, domes, helmets, and lamps, while the third demonstrates sensitivity to surfaces exhibiting scaled or dimpled textures as on golf balls, reptilian skin, and artichokes. Although neuron-level decomposition provides valuable interpretable insights, individual neurons frequently encode multiple concepts simultaneously, or encode concepts incompletely, which complicates conceptual understanding of the neurons and results in more diffuse, uniformly distributed activation patterns in the corresponding heatmaps. Future research should investigate decomposition approaches utilizing linear probing techniques to identify more semantically coherent and interpretable representational subspaces. Alternatively, methods such as PURE (Dreyer et al., 2024) allow us to first disentangle neurons into multiple monosemantic units, which improves the clarity of neuron-level decompositions.

### v3

@@ Appendix H Unifying Concept-Level Explanations, Feature Attribution and Data Attribution
The formulation of DualDA can be combined with concept-level explanations to understand the impact that individual concepts – which may be present or absent in the training and test data – have on their combined DualDA attribution value. For a given concept , assume through a concept-discovery method such as linear probes we have obtained a Concept Activation Vector (CAV) (kim2018interpretability). The larger the dot product between the hidden activations and the vector , the more likely it is that the concept is present in the test sample. We further normalize the CAV . Then, the amount of the DualDA attribution that is explained by concept can be quantified as
@@
An illustrative application of this methodology to an ImageNet sample is presented in Figure H.1. The attribution decomposition is performed with respect to concepts encoded by neurons in the penultimate network layer. Heatmaps are visualized for the three neurons contributing the highest relevance scores. To determine their corresponding conceptual representations, we present the nine training samples that maximize the activation of each neuron: the first neuron responds predominantly to turtle imagery, the second exhibits selectivity for hemispherical structures oriented upward such as shells, domes, helmets, and lamps, while the third demonstrates sensitivity to surfaces exhibiting scaled or dimpled textures as on golf balls, reptilian skin, and artichokes. Although neuron-level decomposition provides valuable interpretable insights, individual neurons frequently encode multiple concepts simultaneously, or encode concepts incompletely, which complicates conceptual understanding of the neurons and results in more diffuse, uniformly distributed activation patterns in the corresponding heatmaps. Future research should investigate decomposition approaches utilizing linear probing techniques to identify more semantically coherent and interpretable representational subspaces. Alternatively, methods such as PURE (dreyer2024pure) allow us to first disentangle neurons into multiple monosemantic units, which improves the clarity of neuron-level decompositions.
Appendix I Preliminary Experiments on Large Language Models
To assess the applicability of our method to Large Language Models, we conduct preliminary experiments that will be expanded in future work. We fine-tune a pre-trained Llama-3.2-1B model (dubey2024llama, released under the Llama 3.1 Community License, available on Huggingface555https://huggingface.co/meta-llama/Llama-3.2-1B) on the AG News dataset for news topic classification (zhang2015character, available on Huggingface666https://huggingface.co/datasets/sh0416/ag_news). The dataset contains 120,000 training samples, each consisting of a news description and a label indicating one of four categories. We assess both the resource consumption for attribution creation and the attribution quality through a Shortcut Detection experiment. All experiments have been run on a NVIDIA H200 Tensor Core GPU.
I.1 Resource Consumption
In this section, we report the computational resources required by DualDA for and compare them to EK-FAC (grosse2023studying), the most computationally efficient variant of Influence Functions available to date. Table I.1 presents the results on the caching time, cache space, average explanation time and VRAM usage for both methods.
Generating attribution scores with EK-FAC for the Llama model demands substantial VRAM. The official implementation777https://github.com/pomonam/kronfluence highlights the parameter per_device_train_batch_size, which allows the user to trade-off VRAM consumption against minor changes in explanation times. To ensure a fair comparison, we report EK-FAC’s results using its most efficient configurations for each metric. Specifically, the VRAM usage reported in Table I.1 corresponds to the absolute minimal setting (i.e. per_device_train_batch_size=1), while the explanation time corresponds to the most efficient setting we obtained with an available VRAM of 145 GB (i.e. per_device_train_batch_size=32).
The precomputation phase of DualDA involves two steps: (1) caching final hidden features and (2) training the SVM surrogate. Feature extraction dominates both caching time and cache size. Table I.1 reports total caching time and cache size, while the training durations for the SVM surrogate models are as follows: 18 minutes 27 seconds for , 48 seconds for , and 32 seconds for . The storage required for the trained SVM parameters requires only 1.9 MB across all values of .
The results show that DualDA achieves improvements of – in caching time, in cache size, at least in average explanation time and at least in VRAM usage.
| DualDA | EK-FAC | Factor of Improvement | |
| Caching Time (s) | 620 – 1695 | 72460 | 42 – 116 |
| Cache Size (GB) | 0.99 | 34.60 | 35 |
| Explanation Time per test sample (s) | 0.04 | ||
| VRAM Usage (GB) | 8.7 |
I.2 Shortcut Detection Experiment
We adapt the task of Shortcut Detection (method:koh_influence; hammoudeh2022identifying), which identifies spurious correlations in training data that enable models to exploit dataset-specific biases, from the visual domain to our language setting. We construct a training set of 4,000 randomly selected samples and introduce a shortcut into 20% of them by randomly reassigning the label to a class cls and inserting the corresponding keyword ‘¡cls¿’ at a random position in the sample text. We apply the same manipulation to create a shortcutted test set consisting of 1,000 test samples. Since models exploiting these shortcuts should rely heavily on the shortcutted training samples, we hypothesize that Data Attribution methods assign high attribution to these samples when explaining predictions on shortcutted test instances. We quantify the overall effect of each training sample on the entire test set by summing its attributions across all shortcutted test instances.
Consistent with our vision domain experiments, lower sparsity improves Shortcut Detection performance. At sparsity levels of or higher, DualDA fails to meaningfully differentiate between shortcutted and non-shortcutted samples. However, at or lower, DualDA successfully identifies shortcuts: it assigns the highest overall attribution to nearly half of all shortcutted samples, and all shortcutted samples are retrieved within the top 50% of ranked training samples. The corresponding Precision-Recall curve (Figure I.1) achieves an AUPRC score of 0.88.